OpenStudy (anonymous):

A van traveling down a slope with a uniform acceleration of 2.15 meters/second2 attains a speed of 20.00 meters/second after 7.00 seconds. What is the initial velocity of the van?

OpenStudy (anonymous):

\[v_{f}= v_{i} + a \Delta t \]

OpenStudy (anonymous):

but they don't give the final velocity

OpenStudy (anonymous):

In fact, they do! The final velocity is the 20 meters per second...

OpenStudy (anonymous):

so would it be 20 divided by acelleration plus seconds

OpenStudy (anonymous):

is the answer 4.95

OpenStudy (johnweldon1993):

The answer is indeed 4.95m/s

OpenStudy (anonymous):

thank you.. could you help me on other questions johnweldon1993

OpenStudy (johnweldon1993):

If I can yes, if not there are others to help :)

OpenStudy (anonymous):

ok thank you hold on

OpenStudy (anonymous):

A car with a mass of 1.1 × 103 kilograms hits a stationary truck with a mass of 2.3 × 103 kilograms from the rear end. The initial velocity of the car is +22.0 meters/second. After the collision the velocity of the car is -11.0 meters/second. What is the velocity of the truck after this elastic collision?

OpenStudy (johnweldon1993):

So for this...It is important to see that momentum is conserved.. meaning \[P_f = P_i\] The momentum before is = to the momentum after so mass of car * Velocity of car(initially) + Mass of truck * velocity of truck(initially) = Mass of car * velocity of car(final) + Mass of truck * velocity of truck(final) Notice you have ALL things...except for Velocity of truck (final)....so rearrange and solve for that :)

OpenStudy (johnweldon1993):

hope that's understandable...not too used to the equation tool lol

OpenStudy (anonymous):

so it would be 22(cars mass) = -11(trucks mass)

OpenStudy (johnweldon1993):

Not sure if you're skipping ahead but.... it would be *rearranged* \[V_truck(final) = \frac{ M_car \times V_car(initial) + M_truck \times V_truck(initial) - M_car \times V_car(final) }{ M_truck }\]

OpenStudy (anonymous):

i don't have a calculator so this is hard

OpenStudy (johnweldon1993):

So the velocity of the truck after the collision would be (1.1 x 10^3kg X 22.0m/s) + (2.3 x 10^3kg X 0) - ( 1.1 x 10^3kg X -11.0m/s) -------------------------------------------------------------------------- 2.3 x 10^3 I don't have a calculator either lol so let me do this on my phone real quick...

OpenStudy (johnweldon1993):

So what I get from doing all this out...is 15.78m/s surprising...I thought it would be less

OpenStudy (johnweldon1993):

Nope it's correct...just checked